The load was connected to the external battery circuit. At a current of 3 A, 18 W of power is
released in the external circuit. Find the load resistance and EMF of the battery if the internal resistance of
the battery is 0.5 .
1
Expert's answer
2020-11-19T06:01:04-0500

Explanations & Calculations


  • The total consumption of power ("\small P" )of the load ("\small R" ) is given by, "\small P = Vi=i^2R=\large \frac{V^2}{R}".
  • Therefore, using the appropriate equation ("\small P=i^2R\,,P=Vi") with the given data (voltage & current), the load resistance & the voltage applied ("\small V_{applied}" ) to the external circuit could be calculated.
  • Therefore,

"\qquad\qquad
\begin{aligned}
\small 18W&= \small (3A)^2\times R_{load}\\
\small R_{load}&= \small \bold{2\Omega}\\
\\
\small V_{applied} &= \small \frac{P}{i}= \frac{18W}{3A}=6V
\end{aligned}"


  • In fact, all the power sources have some internal resistance which drops some voltage across them resulting a lower applied voltage than the EMF of the source.
  • When power is given out of the source to an external circuit, the applied/available voltage is given as "\small V_{applied}=E-i_{applied}R_{internal}".
  • 6V is the applied voltage & the 3A is solely conducted through the source thus the applied current.
  • Therefore, the EMF is,

"\qquad\qquad
\begin{aligned}
\small EMF &= V_{applied}+iR\\
&= \small 6V+(3A\times 0.5\Omega)\\
&= \small \bold{7.5V}
\end{aligned}"